OpenStudy (anonymous):

Of the homes built in Ashe County in one year, 21, or 5 less than 1/3 of all, were of brick. How many homes were built in Ashe County that year?

OpenStudy (anonymous):

1/3x - 5 = 21 and solve right?

OpenStudy (anonymous):

so... 26 = (1/3)x = 78 total?

OpenStudy (anonymous):

thats what I got!!

OpenStudy (anonymous):

:)